OpenStudy (anonymous):

Suppose a bus arrives at a bus stop every 30 minutes. If you arrive at the bus stop at a random time what is the probability that you will have to wait at least 5 min for the bus? (a) 1/6 (b) 1/5 (c) 2/3 (d) 5/6

OpenStudy (anonymous):

C

OpenStudy (kirbykirby):

Let X be the waiting time random variable. Then we can say that X has a continuous uniform distribution, \(X\sim U(0,30)\). Then the question asks to find \(P(X\ge 5)\) \(P(X\ge 5)= 1-P(X \le 5)\) Now recall that the CDF for a conitnuous random variable with distribution \(U(a,b)\) is: \[P(X \ge x)=\frac{x-a}{b-a}, \text{for }x\in[a,b) \] So \[1 - P(X \le 5) =1-\frac{5-0}{30-0}=\frac{5}{6}\]

OpenStudy (anonymous):

thank you

OpenStudy (kirbykirby):

Also, if you think of it intuitively, since at least 5 mins means 5 to 30 mins... that time interval is 25 mins. And 25 min divided by the whole 30 mins = 25/30 = 5/6

OpenStudy (kirbykirby):

yw :)